OpenStudy (anonymous):

Please help me with this mathematical equation? It's not making sense to me in the least. Thank you. Solve the following system using the substitution method. 10x - 3y = -7 5x - y = -2 Answer options: -1/5, -3 1/5,-3 -1/5, 3 1/5, 3

OpenStudy (anonymous):

Take one of the two given equations, substitute the value of y in terms of x then substitute it the other equation.

OpenStudy (anonymous):

how so?

OpenStudy (anonymous):

10x - 3y = -7 5x - y = -2 isolate y from the second : y= 5x + 2 now plug it into the first : 10 -3(5x+2) = -7 now you will get a x value .. can you tell me what is it ?

OpenStudy (anonymous):

OH what am i saying? I understand this

OpenStudy (anonymous):

10 (-1,5) - 3 (-3) = 7

OpenStudy (anonymous):

Answer A is crossed off the list

OpenStudy (anonymous):

Instead of trying to check the answers (which is a valid option in this case) you might try and really solve the system ..

OpenStudy (anonymous):

My error

OpenStudy (anonymous):

Let's continue

OpenStudy (anonymous):

read my first response

OpenStudy (anonymous):

i'm solving for x?

OpenStudy (anonymous):

yes then after we will have x value we can find y value..

OpenStudy (anonymous):

God I don't know this. I seriously don't. I'm alien to this whole stupid concept.

OpenStudy (anonymous):

again : we had two equations : 10x - 3y = -7 5x - y = -2 now, in order to get the solution we have to solve first for one variable and then using the solution for this one variable we will get the other. we can do this by two methods : one- substitution and the other is - elimination in here we have been asked for the substitution : so we find an expression for one of the variables in term of the other from one of the equation in this case i chose to use the second equation and isolate y from it and get : y = 5x + 2 now since i got it from the second equation ill substitute it into the first ill replace y by what i got from the first equation and then ill have an equation with only one variable - x ! which is easy to solve.

OpenStudy (anonymous):

after substitution in the first we have : 10x - 3 * (5x + 2) = -7 tell me please if you understand what i wrote so far

OpenStudy (anonymous):

Why're you multiplying 10x - 3 * (5x + 2) = -7?

OpenStudy (anonymous):

from the second equation i got y = 5x + 2 the original first equation is : 10x - 3y = -7 now i want to substitute the expression of y that i got from the second equation into the first so i replace y by 5x + 2 : 10x - 3 *(5x + 2) = -7

OpenStudy (anonymous):

So by substiting the expression, you multiply? all the time?

OpenStudy (anonymous):

How is that "replacing"?

OpenStudy (anonymous):

the same as if i have x = 3 i know that 4x = 12 since 4 * 3 = 12

OpenStudy (anonymous):

Oh, well that makes sense

terenzreignz (terenzreignz):

Not multiply so much as replacing every instance of 'y' with whatever you came up with (when you isolated y in one of the equations)

OpenStudy (anonymous):

Thank you for clearing that for me :)

OpenStudy (anonymous):

so in this case it is the same i know that y = 5x + 2 then i replace y in the first equation by 5x + 2

OpenStudy (anonymous):

10x - 3 * (5x + 2) = -7 So in order to solve this, we need to find x

OpenStudy (anonymous):

Would I use the second equation to substitute?

OpenStudy (anonymous):

can you solve it now ?: 10x - 3 * (5x + 2) = -7

OpenStudy (anonymous):

this is one equation with one unknown .. just simplify it and isolate x

OpenStudy (anonymous):

It equals -18 without the x's

OpenStudy (anonymous):

10x - 3 *(5x + 2) = -7 10x - [ 3*5x + 3*2 ] = -7 10x - [15x + 6] = -7 10x - 15x - 6 = -7 -5x = -1 x = 1/5

OpenStudy (anonymous):

solving such equation should be familiar to you

OpenStudy (anonymous):

The answer I got is B with my notes

OpenStudy (anonymous):

wait .. now we have x = 1/5 how do you find y ? now you substitute x = 1/5 into one of the original equations.. show me please how you do it

OpenStudy (anonymous):

please read all my notes here and tell me if you understand

OpenStudy (anonymous):

10x - 3y = -75x - y = -2 10(1/5) - 3y

OpenStudy (anonymous):

dont forget the right side of the equation : 10(1/5) - 3y = -7

OpenStudy (anonymous):

so now isolate y .. what do you get ?

OpenStudy (anonymous):

10(1/5) - 3y = -7 To isolate y I would have to substitute? with the other equation?

OpenStudy (anonymous):

no .. this is an equation with only one unknown .. you just simplify and solve for y just the same as i did here : 10x - 3 *(5x + 2) = -7 10x - [ 3*5x + 3*2 ] = -7 10x - [15x + 6] = -7 10x - 15x - 6 = -7 -5x = -1 x = 1/5

OpenStudy (anonymous):

you substitute in order to get an equation with only one variable .. but you have it already so now you just isolate y..

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

so.. solve this: 10(1/5) - 3y = -7

OpenStudy (anonymous):

I understand that I've found x

OpenStudy (anonymous):

To isolate why I subtract something from both sides?

OpenStudy (anonymous):

i dont understand you if you know x = 3 for example tell me what is y if y = 4x ?

OpenStudy (anonymous):

4(3) 12

OpenStudy (anonymous):

and if you knew x = 3 and you wanted to find y if y - 2x = 3 ?

OpenStudy (anonymous):

y - 2(3)= 3 y - 12= 3 12+ 3 = 15 15 - 12 = 3

OpenStudy (anonymous):

where is y :|

OpenStudy (anonymous):

y-2(3)=3 the first line is good now y-6=3 (dont know why you wrote 12) .. now find y

OpenStudy (anonymous):

Huh? Now I'm confused

OpenStudy (anonymous):

you had this : y-2(3)=3 now y - 2*3 = 3 y - 6 = 3 (its not y -12 =3 .. i dont understand why you written -12 instead of -6) now find y using this equation

OpenStudy (anonymous):

How the heck do I solve something I don't know the number to?

OpenStudy (anonymous):

y-6=3 you add 6 to both sides of the equations .. y - 6 + 6 = 3 +6 y = 9 !

OpenStudy (anonymous):

Oh:)

OpenStudy (anonymous):

My final answer's D

OpenStudy (anonymous):

yes because 10(1/5) - 3y = -7 2 - 3y = -7 -3y = -9 y = 3 so 1/5 , 3

OpenStudy (anonymous):

THANK GOD

OpenStudy (anonymous):

please try and solve this system .. i want you to understand and do it : y + x = 3 2y + 3x =7

OpenStudy (anonymous):

show me what you do .. i really want you to understand

OpenStudy (anonymous):

I do, it's all written, step by step :) that's why I said what I did, since I finished it off

OpenStudy (anonymous):

will you try doing this ? y + x = 3 2y + 3x =7 please

OpenStudy (anonymous):

i mean as you wish .. i just want you to be able to solve this by yourself

OpenStudy (anonymous):

One moment

OpenStudy (anonymous):

i want to understand as well, and from what i see on my paper, I do

OpenStudy (anonymous):

ok.. as you wish :)